OpenStudy (anonymous):

Using complete sentences, explain how to completely factor 2x2 − 8x − 42.

jimthompson5910 (jim_thompson5910):

your first task is to factor out the GCF

jimthompson5910 (jim_thompson5910):

what's the GCF in this case?

OpenStudy (anonymous):

2

jimthompson5910 (jim_thompson5910):

factor that out to get what?

OpenStudy (anonymous):

x-4- 21?

jimthompson5910 (jim_thompson5910):

2(x^2 - 4x - 21) is what you should get

jimthompson5910 (jim_thompson5910):

now we must factor x^2 - 4x - 21 how do we do this?

OpenStudy (anonymous):

mulyiply?

jimthompson5910 (jim_thompson5910):

you need to find two numbers that a) multiply to -21 (the last term) AND b) add to -4 (the middle coefficient)

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

what two numbers fit both descriptions?

OpenStudy (anonymous):

7 and 3?

OpenStudy (skullpatrol):

Look for factors of -21 with opposite signs that add up to -4. $$\huge (x+r)(x+s)$$

jimthompson5910 (jim_thompson5910):

close, but notice how 7*3 = positive 21

jimthompson5910 (jim_thompson5910):

one of the numbers has to be negative

OpenStudy (anonymous):

so -7?

jimthompson5910 (jim_thompson5910):

good, -7*3 = -21

OpenStudy (nurali):

2x2 − 8x − 42 2x^2 - 8x - 42 2(x^2-4x-21) 2(x^2-7x+3x-21) 2(x(x-7)+3(x-7)) 2(x - 7)(x + 3)

jimthompson5910 (jim_thompson5910):

and -7 + 3 = -4

jimthompson5910 (jim_thompson5910):

so this means x^2 - 4x - 21 factors to (x-7)(x+3)

jimthompson5910 (jim_thompson5910):

this trick only works if the leading x^2 coefficient is 1

jimthompson5910 (jim_thompson5910):

so the full factorization is 2(x-7)(x+3)

OpenStudy (anonymous):

ok thank you!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (skullpatrol):

Check the result by multiplying the factors.

OpenStudy (anonymous):

when the product of a & c is negative always subtract a number from a bigger number to get b.i.e., p-q=b